OpenStudy (anonymous):

Factor (3x-7)(x+2)=0

OpenStudy (anonymous):

It's already factored isn't it ?

OpenStudy (anonymous):

oh sorry i mean solve

OpenStudy (mimi_x3):

x = 7/3 , -2

OpenStudy (anonymous):

where are you stuck here ?

OpenStudy (anonymous):

do you know that: (3x-7)(x+2)=0 means that either (3x-7)=0 or: (x+2)=0

OpenStudy (anonymous):

oh i get you man

OpenStudy (anonymous):

so how would you solve, now that you know this ?

OpenStudy (mimi_x3):

I already gave him the answer xD

OpenStudy (anonymous):

yeah - but wanted to check if he understands how to do it in the future.

OpenStudy (mimi_x3):

lols , its common sense anyway xD

OpenStudy (anonymous):

Id plug in 0 in to either(3x-7)(x+2) and just you know math it up from thereee

OpenStudy (anonymous):

after you do it yourself the first time, it becomes common sense :)

OpenStudy (mimi_x3):

no , you don't plug in the zero

OpenStudy (mimi_x3):

your not finding the intercept or anything xD